answer to 5(4-7u)=-u-48
Answer by DrBeeee(684)

Given:
(1) 5*(4-7u) = -u - 48
Your first step is (usually always) to evaluate the parentheses.
Then (1) becomes
20 - 35u = -u - 48
Combine like terms yields
68 = 34u or
34u = 68
Now divide each side by 34 and obtain
34u/34 = 68/34 or
(2) u = 2
ALWAYS check your answer by substituting your answer into the ORIGINAL equation, (1) in this case, to see if you get equality.
Is [5(4 - 7*2) = -2 -48]?
Is [(20 - 5*14} = -50]?
Is [ 20 - 70) = -50]?
Is [-50 = -50]? Yes
Answer: The solution is u = 2.